40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 15 11:38:31 2009, rob, Update, Locking, MC_F channel dead mcfdead.png
    Reply  Thu Jul 16 14:41:38 2009, rob, Update, Locking, MC_F channel dead IOO_ICS_0_15.pngIOO_ICS_15_32.png
       Reply  Thu Jul 16 14:54:05 2009, rob, Update, Locking, MC_F channel dead 
Message ID: 1758     Entry time: Thu Jul 16 14:41:38 2009     In reply to: 1747     Reply to this: 1759
Author: rob 
Type: Update 
Category: Locking 
Subject: MC_F channel dead 

Quote:

It's railed.  This is what halted locking progess on Monday night, as this channel is used for the offloadMCF script, which slowly feeds back a CARM signal to the ETMs to prevent the VCO from saturating.

 

Attached is a 5 day trend, which shows that the channel went dead a few days ago.  All the channels shown are being collected from the same ICS110B (I think), but only some are dead.  It looks like they went dead around the time of the "All computers down" from Sunday.

 Attached are the channels being recorded from the ICS110B in 1Y2 (the IOO rack).  Channels 12, 13, 16, 17, 22, 24, 25 appear to have gone dead after the computer problems on Sunday.

Attachment 1: IOO_ICS_0_15.png  25 kB  | Hide | Hide all
IOO_ICS_0_15.png
Attachment 2: IOO_ICS_15_32.png  22 kB  | Hide | Hide all
IOO_ICS_15_32.png
ELOG V3.1.3-